Shelly can row 5mp in still water. It takes her as long to row 7miles upstream as it does 23 miles downstream. How fast is the current?
Answer by Alan3354(30993)

Shelly can row 5mp in still water. It takes her as long to row 7miles upstream as it does 23 miles downstream. How fast is the current?
---------
c = speed
t = 7/(5-c) = 23/(5+c)
7(5+c) = 23(5-c)
35 + 7c = 115 -23c
30c = 80
c = 8/3 mi/hr